# Additive Models Explained: A Computational Complexity Approach

### v1

@@ 6.2 Pseudo Polynomial-Time Explanation Computations
Table 1 shows that both CC and SHAP-C queries are intractable. However, if the GAM’s weights $\beta_{0},\beta_{1},\ldots,\beta_{n}$ are given in unary (rather than binary), they become computable in polynomial time. This pseudo-polynomial setting enables efficient computation when weight precision is bounded, using dynamic programming similar to (BaMoPeSu20, ), and detailed in our proofs. It applies to enumerable discrete domains (all component types), where enumeration makes $\mathbb{E}_{\textbf{z}\sim\mathcal{D}_{p}}[f(\textbf{x}_{i})]$ tractable.
@@
Given any GAM over an enumerable discrete input, the CC and SHAP-C queries can be computed in pseudo-polynomial time.
@@ 6.3 Intractable Explanation Computations
Let there be a NAM or an EBM over a discrete or continuous input space, then solving the CSR, and MSR queries are coNP-Complete, and the MCR query is NP-Complete.
@@
Intractability of SHAP and count completions. Finally, we show that computing the SHAP and CC queries is generally #\#P-Hard by reducing to the model counting problem (van2022tractability, ), implying intractability for EBMs and NAMs. While other settings are also intractable (see Subsection 6.2), their complexity can be mitigated using pseudo-polynomial algorithms:
@@ 7 Practical Implications
    Input domain transformations. We find that the computational complexity of generating explanations is highly sensitive to the structure of the input domain. Small changes in the domain can shift problems from intractable to tractable. This insight motivates the use of input transformations, such as discretization, a standard step in many GAM pipelines (caruana2015intelligible, ; nori2021accuracy, ), to reshape the input space in ways that make explanation computation efficient.
@@
    Quantization. Our pseudo-polynomial algorithms show that explanation problems that are intractable in full generality become tractable when GAM component weights are quantized. This suggests a promising practical approach: by quantizing model coefficients, one can enable efficient computation of explanations that would otherwise be infeasible.
@@ 8 Related Work
Our work builds upon prior research exploring the computational complexity of obtaining various types of explanations for ML models (BaMoPeSu20, ; arenas2022computing, ; blanc2021provably, ; adolfi2024computational, ; blanc2022query, ; huang2023feature, ; barcelo2025explaining, ; bhattacharjee2024auditing, ; kozachinskiy2023inapproximability, ; marzouk2024tractability, ; laber2024computational, ; bounia2024enhancing, ; xue2023stability, ; jin2025probabilistic, ). This area is closely tied to the subfield of interest known as formal explainable AI (marques2022delivering, ; bassan2023towards, ) which focuses on explanations with mathematical guarantees (ignatiev2019abduction, ; wu2024verix, ; izza2021efficient, ; audemard2020tractable, ; darwiche2020reasons, ; bassanfast, ; izza2023locally, ; bassan2023formally, ; bouniausing, ), where analyzing the complexity of producing explanations with such guarantees plays a central role (izza2023computing, ; ordyniak2023parameterized, ; audemard2021computational, ; cooper2021tractability, ; arenas2022computing, ; ordyniak2024explaining, ; van2022tractability, ).
Previous work has mainly examined the complexity of generating explanations for specific ML models, including black-box models like neural networks (adolfi2024complexity, ; BaMoPeSu20, ) and interpretable models like decision trees (arenas2022computing, ; bounia2023approximating, ; huang2021efficiently, ). Closer to our work are studies on linear models (BaMoPeSu20, ; subercaseaux2024probabilistic, ; marques2020explaining, ) — a simplified form of GAMs — typically under binary input-output settings. In contrast, we offer a comprehensive analysis of a broad class of GAMs, covering diverse component models, explanation types, input assumptions, and approximation results. We elaborate further on related works in Appendix A.
Lastly, we note that some terms in our work have sometimes been referred to differently in prior literature. For instance, while the term sufficient reasons is widely used (BaMoPeSu20, ; bassan2025self, ; darwiche2020reasons, ), it is also referred to as abductive explanations (ignatiev2019relating, ). Moreover, subset-minimal sufficient reasons are closely related — though not identical — to prime implicants in Boolean formulas (darwiche2002knowledge, ; darwiche2020reasons, ; shih2018symbolic, ).Similarly, the CC query corresponds to the $\delta$-relevant set waldchen2021computational ; izza2021efficient ; subercaseaux2024probabilistic , which checks if the completion count exceeds a given threshold $\delta$.
@@ 9 Limitations and Future Work
More concretely, our results open the door to a wide range of future practical implementations of our algorithms (see Section 9). From a theoretical perspective, our analysis can be extended in several directions. A first promising avenue is to better understand how high-order interactions, which are widely used in GAMs (changnode, ; lou2013accurate, ; kim2022higher, ; enouen2025instashap, ), affect the complexity of generating explanations, and how this varies across explanation types and component models. Another interesting direction is to study how structural assumptions over the GAM influence complexity. For instance, one may consider the effect of concurvity (kovacs2024feature, ; siems2023curve, ). In the extreme case of high concurvity, where smooth components are strongly correlated and behave almost like a single spline, the GAM may resemble a simpler function class, potentially making some explanation problems easier. Conversely, in the case of low concurvity, especially for explanation types that rely on interactions, the high separability between components may also offer computational advantages. Finally, for intractable settings, one can attempt to leverage automated reasoning tools, such as neural network verifiers (wu2024marabou, ; wang2021beta, ) or MILP/MaxSAT solvers for tree ensembles (ignatiev2022using, ; chen2019robustness, ) to obtain explanations. Moreover, while we propose two central strategies for circumventing intractability, input transformations and quantization (see Section 9), exploring additional circumvention approaches such as approximation guarantees, probabilistic or statistical relaxations, and PAC-based guarantees remains an important direction for future research.
@@ Appendix H Proof of Proposition 4
Given any GAM over an enumerable discrete input, or a Smooth GAM over a discrete or continuous input, the CC and SHAP-C queries can be obtained in pseudo-polynomial time.
@@
Given any GAM over an enumerable discrete input setting, the CC query can be solved in pseudo-polynomial time
@@
We observe that this problem resembles the classic #\#Knapsack problem—the counting variant of Knapsack — except that each variable $x_{i}$ ranges over a set $S_{i}$ of enumerable integers, rather than being restricted to $\{0,1\}$. The work of [16] reduces the CC query for Perceptrons (i.e., linear models over binary inputs with unary weights) to the classic #\#Knapsack problem via a polynomial-time reduction, where $C$ is set to the bias $b$ of the linear model. We observe that applying the same reduction — except replacing binary inputs with sets of enumerable discrete values — yields an instance of the #\#Multi-Choice Knapsack problem instead (and the target integer $C$ still equals the bias $b$ of the linear model. Thus, it remains to show that #\#Multi-Choice Knapsack is solvable in pseudo-polynomial time, which we do by adapting the standard dynamic programming algorithm typically used for this problem (and also discussed in [16]).
@@
Given any GAM over an enumerable discrete input setting, the SHAP query for classification tasks can be solved in pseudo-polynomial time.
@@ Appendix I Proof of Proposition 5
Given $f$ we construct $f^{\prime}$ such that the first input layer of $f^{\prime}$ becomes the second hidden layer of $f$ and we add an additional input layer for $f^{\prime}$ consisting of only one input, as well as an additional hidden layer (which is the first hidden layer of $f^{\prime}$, that is twice as larger than that of the input layer of $f$. We connect the single input layer of $f^{\prime}$ to all neurons in the first hidden layer. Each pair of neurons in the first layer is connected to a single hidden layer in the second hidden layer (which corresponds to the original input of $f$). We set the biases of all of the neurons in both the first and second hidden layers. Let us assume some neuron $x^{2}_{i}$ at the second hidden layer which represents $\textbf{x}_{i}$ from the original input $f$. Let $x^{1}_{i}$ and $x^{2}_{i}$ denote the pair of two neurons that are connected to it from the first layer of $f^{\prime}$. We set the weight that corresponds with
@@
We observe that, as previously mentioned, the proof in [96] was derived from the CNF-SAT problem, initially assuming the input space $\mathcal{X}:={0,1}^{n}$. Consequently, we will specifically demonstrate hardness for the neural network verification problem involving a model of the form $f^{\prime}:\mathbb{R}\to\mathbb{R}$ by reducing it from the neural network verification problem over a model of the form $f:{0,1}^{n}\to\mathbb{R}$.
@@
Reducing the continuous domain $0,2^{n}-1]$ for an individual feature to the discrete domain $\{0,1\}^{n}$ for $n$ features. The construction of $f^{\prime}$ is as follows. Following the first input neuron $o^{0}$ we will connect it to a neuron in the preceding layers which calculates: $o^{1}_{0}:=\text{ReLU}(o^{0}-2^{n-1})$ and $o^{1}_{1}:=|o^{0}-2^{n-1}|$. Where $o^{1}_{1}:=|o^{0}-2^{n-1}|$ can be computed given the explanation from before. We connect $o^{1}_{0}$ to the first neuron from layer $o^{\prime}_{1}$ via a linear transformation, i.e., $o^{\prime}_{1}:=o^{1}_{0}$.
We now move on and construct another hidden layer over which we compute $o^{2}_{0}:=\text{ReLU}(o^{1}_{1}-2^{n-2})$ and $o^{2}_{1}:=|o^{1}_{1}-2^{n-2}|$. We connect $o^{2}_{0}$ to the second neuron in the $o^{\prime}$ layer via a linear transformation, i.e., we compute $o^{\prime}_{2}:=o^{2}_{0}$. We then continue performing this entire process recursively, i.e., for the $i$’th iteration in this process we construct $o^{i}_{0}:=\text{ReLU}(o^{i-1}_{1}-2^{n-i}$, $o^{i}_{1}:=|o^{i-1}_{1}-2^{n-i}|$, and connect input $i$ in the $o^{\prime}$ layer to $o^{i}_{0}$ via a linear transformation, i.e.,: $o^{\prime i}:=o^{i}_{0}$.
The following construction takes some arbitrary assignment $o_{0}$ in the range $0,2^{n}]$ and performs a set of binary decisions, each time decreasing $2^{i-1}$ and continuing with the remainder. This process hence gives us in the first neuron $o^{\prime}_{1}$ a positive value iff $o_{0}$ is larger than $2^{n-1}$, and otherwise will be assigned a 0. The second neuron $o^{\prime}_{2}$ will get a positive value iff $|o^{\prime}-2^{n-1}|-2^{n-2}$ is larger than 0, and otherwise it will be set to 0. The third neuron $o^{\prime}_{3}$ will get a positive value iff $||o^{\prime}-2^{n-1}|-2^{n-2}|-2^{n-3}$ is larger than 0, and otherwise, it will be assigned 0. This continues for every $i$. It is straightforward to show that for a value $o^{\prime}=\lfloor x\rfloor$ for some $0\geq x\geq 2^{n}$ it holds that we can equivalentley describe $o^{\prime}$ as a binary vector of size $n$ which is equivalent to the binary value of $x$, where each “1” value in place $i$ in the binary vector corresponds to some positive assignment in place $i$ in the vector $o^{\prime}$, and each “0” value in place $i$ in the binary vector corresponds to some 0 assignment in the $o^{\prime}$ vector.
@@
Proof. Membership. We will start by establishing the claim for the general discrete setting. Since each domain $\mathcal{X}_{i}$ for feature $\textbf{x}_{i}$ is defined by a minimum and maximum value represented as a set of $q$ binary features, and since each value in the domain consists of at most $q$ binary features, we can non deterministically guess input assignments for each feature (corresponding to guessing the values of the binary encoding). Consequently, we obtain a value $\textbf{z}\in\mathbb{F}$ based on our guess. Analogous to the process outlined in both algorithms 1 and 2, we determine the “penalty” value for each feature $i$, but now this is done not for the entire domain $\mathcal{X}_{i}$, but solely with respect to the guessed input $\textbf{z}_{i}$. Specifically, we define:
$p_{i}:=\begin{cases}\text{min}\{\beta_{i}\cdot\textbf{x}_{i}\ ,\ \beta_{i}\cdot\textbf{z}_{i}\}\quad if\ \ f(\textbf{x})>0\\ \text{max}\{\beta_{i}\cdot\textbf{x}_{i}\ ,\ \beta_{i}\cdot\textbf{z}_{i}\}\quad 0\quad otherwise\end{cases}$ (30)
Furthermore, similarly to algorithms 1 and 2, we will require a measure for the value of feature $i$ when it is fixed. We define $v_{i}:=\beta_{i}\cdot\textbf{x}_{i}-p_{i}$. Note that when $\sum_{i\in[n]}\beta_{i}\cdot\textbf{x}_{i}+\beta_{0}<0$, i.e., when $f(\textbf{x})=0$, a “good score” will be a highly negative one. Thus, we specifically define $v_{i}:=(\beta_{i}\cdot\textbf{x}_{i}-p_{i})\cdot\text{sign}(\sum_{i\in[n]}\beta_{i}\cdot\textbf{x}_{i}+\beta_{0})$. However, for simplicity, we will assume without loss of generality that $f(\textbf{x})>0$, and therefore $v_{i}:=\beta_{i}\cdot\textbf{x}_{i}-p_{i}$. Intuitively, $v_{i}$ reflects the value of feature $i$ when it is held constant rather than allowed to take on any value, thereby influencing the penalty score $p_{i}$.
We proceed with the following procedure: first, we sort all features $i\in[n]$ in descending order based on their value $v_{i}:=\beta_{i}\cdot x_{i}-p_{i}$. Next, we define $S$ as the set of the last nkn-k features in this ordering (i.e., the nkn-k features with the lowest ranking). Finally, we compute:
$\text{sign}(\sum_{i\in\bar{S}}\beta_{i}\cdot x_{i}+\sum_{j\in S}p_{j}+\beta_{0}\big)\cdot\text{sign}(\sum_{i\in n}\beta_{i}\cdot x_{i}+\beta_{0})<0$ (31)
And if this condition holds, the algorithm concludes that there is no sufficient reason of size $k$ or less for $\langle f,\textbf{x}\rangle$ (and this will determine membership in coNP). We will now explain the correctness of the procedure: we have demonstrated in algorithms 1 that $\text{sign}(\sum_{i\in\bar{S}}\beta_{i}\cdot x_{i}+\sum_{j\in S}p_{j}+\beta_{0})\cdot\text{sign}(\sum_{i\in n}\beta_{i}\cdot x_{i}+\beta_{0})<0$ is true if and only if $S$ is a contrastive reason of $\langle f,\textbf{x}\rangle$ (equivalently, this means $\overline{S}$ is not a sufficient reason). Recall that we have sorted the values in descending order of $v_{i}:=\beta_{i}\cdot\textbf{x}_{i}-p_{i}$, which determines that the subset $S$ is the subset of size nkn-k that minimizes the difference:
$(\sum_{i\in n}\beta_{i}\cdot x_{i}+\beta_{0})-(\sum_{i\in\bar{S}}\beta_{i}\cdot x_{i}+\sum_{j\in S}p_{j}+\beta_{0}\big)=\sum_{j\in S}(\beta_{j}\cdot\textbf{x}_{j}-p_{j})$ (32)
This implies that, since for $S$ it holds that $\text{sign}(\sum_{i\in\bar{S}}\beta_{i}\cdot x_{i}+\sum_{j\in S}p_{j}+\beta_{0}\big)\cdot\text{sign}(\sum_{i\in n}\beta_{i}\cdot x_{i}+\beta_{0})<0$, it must necessarily follow that for any other subset $S^{\prime}$ of size nkn-k, $\text{sign}(\sum_{i\in\bar{S^{\prime}}}\beta_{i}\cdot x_{i}+\sum_{j\in S^{\prime}}p_{j}+\beta_{0}\big)\cdot\text{sign}(\sum_{i\in n}\beta_{i}\cdot x_{i}+\beta_{0})<0$. This establishes that $S^{\prime}$ is contrastive. Consequently, every subset of size nkn-k is a contrastive reason for $\langle f,\textbf{x}\rangle$, which further implies that any subset of size nkn-k or larger serves as a contrastive reason for $\langle f,\textbf{x}\rangle$. In summary, this demonstrates that no sufficient reason of size $k$ or less exists for $\langle f,\textbf{x}\rangle$, which demonstrates membership in coNP.
For the continuous setting, we first observe that determining whether a satisfying assignment exists over a continuous input domain for a ReLU-based neural network, given piecewise-linear input and output specifications, can be formulated as a piecewise-linear program, as discussed in prior works [96, 69]. Specifically, when the activation status of each ReLU in the neural network is fixed as either active or inactive, the problem of verifying whether a set of linear constraints on the input (e.g., upper and lower bounds) can produce a specified output (e.g., ensuring the target class exceeds all others) reduces to solving a linear program. The "exponential" complexity arises from enumerating all possible activation combinations, as each ReLU can either activate or remain inactive. Indeed, the proof provided by [96] to establish NP membership for this problem involves guessing the ReLU activations and solving the corresponding linear program. However, the intricate structure of the MSR query in our case prevents directly applying such an approach, necessitating a different solution method.
We observe that for a set of $m$ constraints in a linear program, each represented as a rational number $\frac{p_{i}}{q_{i}}$, it is straightforward to demonstrate that the linear program can be solved within a tolerance of $\frac{1}{q_{1}\cdot q_{2}\cdot\ldots\cdot q_{m}}$. Typically, tighter tolerances are achieved by considering subdeterminants [42] or by assuming approximations in the input. Since determining the existence of a satisfying assignment corresponds to a piecewise-linear program, the tolerance for solving this problem must account for each ReLU activation (either activated or inactivated), resulting in the smallest possible tolerance overall. It is evident that for any given constraint (including those on the input, weights, and biases, all of which involve inequalities or summations over terms of the form $\frac{p_{i}}{q_{i}}$), selecting the largest $q_{i}$ from all $m$ constraints (let us denote it by $q^{*}$) shows that the tolerance is bounded by $\frac{1}{(q^{*})^{m}}$.
Naturally, the binary representation of inputs within this tolerance can be encoded using $m\cdot\log(q^{*})$ bits, which remains polynomial. By enumerating all possible assignments under this precision, it is possible to determine whether a satisfying output assignment exists for any given domain over the inputs (defined by linear constraints). Thus, we can apply the exact procedure described earlier for the general discrete setting, which involves guessing assignments (equivalent to guessing the bit values of each input) by extending it to this instance. Specifically, given the model $f$, we can define a piecewise linear program that represents the verification query over $f$ as outlined in [96], incorporating upper and lower bounds on the input and output. From there, we calculate an upper bound for the tolerance and ensure the number of bits required to validate the existence of a satisfying assignment for $f$ is maintained (using $m\cdot\log(q^{*})$). We then proceed with the same approach described for the general discrete setting to demonstrate membership within coNP.
@@
Hardness. We have proven in Lemma 14 that the neural network verification problem for a neural network with one input and one output, where the input domain is either continuous or general discrete, is NP-Hard. We can now leverage this hardness result to demonstrate that the MSR query for a NAM defined over either a general discrete or a continuous domain is coNP-Hard.
@@
This establishes that obtaining the MSR query for a NAM, where $\mathcal{X}$ is defined as either a general discrete or a continuous input domain, is coNP-Hard.

### v2

@@ 6.2 Pseudo Polynomial-Time Explanation Computations
Table 1 shows that both CC and SHAP-C queries are intractable. However, if the GAM’s internal encoding and input is given in unary (rather than binary) and the GAM has a polynomially bounded output, they become computable in polynomial time. This pseudo-polynomial setting enables efficient computation when the encoding precision is bounded, using dynamic programming similar to (Barceló et al., 2020), and detailed in our proofs. It applies to enumerable discrete domains (all component types), where enumeration makes $\mathbb{E}_{\textbf{z}\sim\mathcal{D}_{p}}[f(\textbf{x}_{i})]$ tractable.
@@
Given any unary-encoded GAM over a discrete, enumerable input domain with polynomially bounded output, the CC and SHAP-C queries admit a pseudo-polynomial-time algorithm.
@@ 6.3 Intractable Explanation Computations
Let there be a NAM or an EBM over a discrete or continuous input space, then solving the CSR query is coNP-Complete, the MSR query is coNP-Hard, and the MCR query is NP-Complete.
@@
Intractability of SHAP and count completions. Finally, we show that computing the SHAP and CC queries is generally #\#P-Hard by reducing to the model counting problem (Van den Broeck et al., 2022), implying intractability for EBMs and NAMs. While other settings are also intractable (see Subsection 6.2), their complexity can be mitigated using pseudo-polynomial algorithms:
@@ 7 Practical Implications
    Input domain transformations. We find that the computational complexity of generating explanations is highly sensitive to the structure of the input domain. Small changes in the domain can shift problems from intractable to tractable. This insight motivates the use of input transformations, such as discretization, a standard step in many GAM pipelines (Caruana et al., 2015; Nori et al., 2021), to reshape the input space in ways that make explanation computation efficient.
@@
    Quantization. Our pseudo-polynomial algorithms show that explanation problems that are intractable in full generality become tractable when the GAM encoding is quantized. This suggests a promising practical approach: by quantizing model coefficients, one can enable efficient computation of explanations that would otherwise be infeasible.
@@ 8 Related Work
Our work builds upon prior research exploring the computational complexity of obtaining various types of explanations for ML models (Barceló et al., 2020; Arenas et al., 2022; Blanc et al., 2021; Adolfi et al., 2025; Blanc et al., 2022; Huang et al., 2023; Barceló et al., 2025; Bhattacharjee and Luxburg, 2024; Kozachinskiy, 2023; Marzouk and De La Higuera, 2024; Laber, 2024; Bounia, 2024; Xue et al., 2023; Jin et al., 2025). This area is closely tied to the subfield of interest known as formal explainable AI (Marques-Silva and Ignatiev, 2022; Bassan and Katz, 2023) which focuses on explanations with mathematical guarantees (Ignatiev et al., 2019b; Wu et al., 2023; Izza et al., 2021; Audemard et al., 2020; Darwiche and Hirth, 2020; Bassan et al., 2025a; Izza et al., 2023b; Bassan et al., 2023; Bounia, 2025), where analyzing the complexity of producing explanations with such guarantees plays a central role (Izza et al., 2023a; Ordyniak et al., 2023; Audemard et al., 2021; Cooper and Marques-Silva, 2021; Arenas et al., 2022; Ordyniak et al., 2024; Van den Broeck et al., 2022).
Previous work has mainly examined the complexity of generating explanations for specific ML models, including black-box models like neural networks (Adolfi et al., 2024; Barceló et al., 2020) and interpretable models like decision trees (Arenas et al., 2022; Bounia and Koriche, 2023; Huang et al., 2021). Closer to our work are studies on linear models (Barceló et al., 2020; Subercaseaux et al., 2025; Marques-Silva et al., 2020) — a simplified form of GAMs — typically under binary input-output settings. In contrast, we offer a comprehensive analysis of a broad class of GAMs, covering diverse component models, explanation types, input assumptions, and approximation results. We elaborate further on related works in Appendix A.
Lastly, we note that some terms in our work have sometimes been referred to differently in prior literature. For instance, while the term sufficient reasons is widely used (Barceló et al., 2020; Bassan et al., 2025d; Darwiche and Hirth, 2020), it is also referred to as abductive explanations (Ignatiev et al., 2019a). Moreover, subset-minimal sufficient reasons are closely related — though not identical — to prime implicants in Boolean formulas (Darwiche and Marquis, 2002; Darwiche and Hirth, 2020; Shih et al., 2018).Similarly, the CC query corresponds to the $\delta$-relevant set Wäldchen et al. (2021); Izza et al. (2021); Subercaseaux et al. (2025), which checks if the completion count exceeds a given threshold $\delta$.
@@ 9 Limitations and Future Work
More concretely, our results open the door to a wide range of future practical implementations of our algorithms (see Section 9). From a theoretical perspective, our analysis can be extended in several directions. A first promising avenue is to better understand how high-order interactions, which are widely used in GAMs (Chang et al., 2022; Lou et al., 2013; Kim et al., 2022; Enouen and Liu, 2025), affect the complexity of generating explanations, and how this varies across explanation types and component models. Another interesting direction is to study how structural assumptions over the GAM influence complexity. For instance, one may consider the effect of concurvity (Kovács, 2024; Siems et al., 2023). In the extreme case of high concurvity, where smooth components are strongly correlated and behave almost like a single spline, the GAM may resemble a simpler function class, potentially making some explanation problems easier. Conversely, in the case of low concurvity, especially for explanation types that rely on interactions, the high separability between components may also offer computational advantages. Finally, for intractable settings, one can attempt to leverage automated reasoning tools, such as neural network verifiers (Wu et al., 2024; Wang et al., 2021b) or MILP/MaxSAT solvers for tree ensembles (Ignatiev et al., 2022; Chen et al., 2019) to obtain explanations. Moreover, while we propose two central strategies for circumventing intractability, input transformations and quantization (see Section 9), exploring additional circumvention approaches such as approximation guarantees, probabilistic or statistical relaxations, and PAC-based guarantees remains an important direction for future research.
@@ Appendix H Proof of Proposition 4
Given any unary-encoded GAM over a discrete, enumerable input domain with polynomially bounded output, the CC and SHAP-C queries admit a pseudo-polynomial-time algorithm.
@@
Given any unary-encoded GAM over a discrete, enumerable input domain with polynomially bounded output, the CC query admits a pseudo-polynomial-time algorithm.
@@
We observe that this problem resembles the classic #\#Knapsack problem — the counting variant of Knapsack — except that each variable $x_{i}$ ranges over a set $S_{i}$ of enumerable integers, rather than being restricted to $\{0,1\}$. This range of possible values corresponds to the outputs of the GAM. Since we assume that the GAM uses an enumerably discrete unary input encoding and that its outputs are polynomially bounded, it follows that the set of possible outputs is itself polynomially bounded with respect to the input encoding (and therefore also unary). The work of [15] reduces the CC query for Perceptrons (i.e., linear models over binary inputs with unary weights) to the classic #\#Knapsack problem via a polynomial-time reduction, where $C$ is set to the bias $b$ of the linear model. We observe that applying the same reduction — except replacing binary inputs with sets of enumerable discrete values representing the enumerable set of unary-encoded possible outputs of the GAM — yields an instance of the #\#Multi-Choice Knapsack problem instead (and the target integer $C$ equals the intercept $\beta_{0}$ of the GAM instead of the bias $b$ of the linear model). Thus, it remains to show that #\#Multi-Choice Knapsack is solvable in pseudo-polynomial time, which we do by adapting the standard dynamic programming algorithm typically used for this problem (and also discussed in [15]).
@@
Given any unary-encoded GAM over a discrete, enumerable input domain with polynomially bounded output, the SHAP query admits a pseudo-polynomial-time algorithm.
@@ Appendix I Proof of Proposition 5
We observe that, as previously mentioned, the proof in [95] was derived from the CNF-SAT problem, initially assuming the input space $\mathcal{X}:=\{0,1\}^{n}$. Consequently, we will specifically demonstrate hardness for the neural network verification problem involving a model of the form $f^{\prime}:\mathbb{R}\to\mathbb{R}$ by reducing it from the neural network verification problem over a model of the form $f:\{0,1\}^{n}\to\mathbb{R}$.
@@
Reducing the continuous domain $0,2^{n}-1]$ for an individual feature to the discrete domain $\{0,1\}^{n}$ for $n$ features. The construction of $f^{\prime}$ is as follows. Following the first input neuron $o^{0}$ we will connect it to a neuron in the preceding layers which calculates: $o^{1}_{0}:=\text{ReLU}(o^{0}-2^{n-1})$ and $o^{1}_{1}:=\text{ReLU}(o^{1}_{0})-\text{ReLU}(o^{1}_{0}-2^{n-1})$. We connect $o^{1}_{1}$ to the first neuron from layer $o^{\prime}_{1}$ via a linear transformation, i.e., $o^{\prime}_{1}:=o^{1}_{1}$. We also set $o^{1}_{2}:=o^{1}_{0}-\text{ReLU}(o^{1}_{1})$.
We now move on and construct another hidden layer over which we compute $o^{2}_{0}:=\text{ReLU}(o^{1}_{1}-2^{n-2})$ and $o^{2}_{1}:=\text{ReLU}(o^{1}_{1})-\text{ReLU}(o^{1}_{1}-2^{n-2})$. We connect $o^{2}_{1}$ to the second neuron in the $o^{\prime}$ layer via a linear transformation, i.e., we compute $o^{\prime}_{2}:=o^{2}_{1}$. We then again set $o^{2}_{2}:=o^{2}_{0}-\text{ReLU}(o^{2}_{1})$ and $o^{3}_{0}:=\text{ReLU}(o^{2}_{1}-2^{n-3})$, and continue performing this entire process recursively, i.e., for the $i$’th iteration in this process we construct $o^{i}_{0}:=\text{ReLU}(o^{i-1}_{2}-2^{n-i})$, $o^{i}_{1}:=\text{ReLU}(o^{i}_{1})-\text{ReLU}(o^{i}_{0}-2^{n-i})$$o^{i}_{2}:=o^{i}_{0}-\text{ReLU}(o^{i}_{1})$ and connect input $i$ in the $o^{\prime}$ layer to $o^{i}_{1}$ via a linear transformation, i.e.,: $o^{\prime i}:=o^{i}_{1}$.
The following construction takes some arbitrary assignment $o_{0}$ in the range $0,2^{n}]$ and performs a set of binary decisions, each time decreasing $2^{i-1}$ and continuing with the remainder. This process hence gives us in the first neuron $o^{\prime}_{1}$ a positive value iff $o_{0}$ is larger than $2^{n-1}$, and otherwise will be assigned a 0. The second neuron $o^{\prime}_{2}$ will get a positive value iff the remainder after decreasing $2^{n-1}$ is larger than 0, and otherwise it will be set to 0. The third neuron $o^{\prime}_{3}$ will get a positive value iff the remainder of what is left, while decreasing $2^{n-2}$ from it is larger than 0, and otherwise, it will be assigned 0. This continues for every $i$. It is straightforward to show that for a value $o^{\prime}=\lfloor x\rfloor$ for some $0\geq x\geq 2^{n}$ it holds that we can equivalentley describe $o^{\prime}$ as a binary vector of size $n$ which is equivalent to the binary value of $x$, where each “1” value in place $i$ in the binary vector corresponds to some positive assignment in place $i$ in the vector $o^{\prime}$, and each “0” value in place $i$ in the binary vector corresponds to some 0 assignment in the $o^{\prime}$ vector.
@@
Proof. Hardness. We have proven in Lemma 14 that the neural network verification problem for a neural network with one input and one output, where the input domain is either continuous or general discrete, is NP-Hard. We can now leverage this hardness result to demonstrate that the MSR query for a NAM defined over either a general discrete or a continuous domain is coNP-Hard.
@@
This establishes that obtaining the MSR query for a NAM, where $\mathcal{X}$ is defined as either a general discrete or a continuous input domain, is coNP-Hard. We note that a recent result by [22] strengthens this bound for the optimization variant, showing that the continuous version of the problem is in fact OptP[$\mathcal{O}(\log n)$]-hard under metric reductions (see Section B.5 of the appendix).
Regarding membership, although showing that the problem lies in $\Sigma_{2}^{P}$ is straightforward (as the same task is known to be in $\Sigma_{2}^{P}$ for general neural networks [15]), we refer the reader to a recent result in [22], which shows that the optimization variant of the problem is in $\mathrm{FP}^{\mathrm{NP}}[\mathcal{O}(n)]$ (see Section B.5 of the appendix). This, in turn, implies that the corresponding decision variant considered here belongs to $\Delta_{2}^{P}$.